could you help me to solve the equation 6y−9=4x+9 for y?
I found that it has more than one answer, but the website doesn't accept.
Thank you.

Answer by Alan3354(69443)

6y - 9 = 4x + 9
6y = 4x + 18
y = (4x + 18)/6 = (2x + 9)/3
or y = (2/3)x + 3